# Multimodal Fusion Balancing Through Game-Theoretic Regularization

### v2

@@ A.5 MCR on M modalities
$I(X_{1},..,X_{M};Y)=\sum_{\mathcal{S}\subseteq\{X_{1},..,X_{M}\},\mathcal{S}\neq\emptyset}I(\mathcal{S};Y\mid\{X_{1},\dots,X_{M}\}\setminus\mathcal{S})+I(X_{1},..,X_{M})-I(X_{1},..,X_{M}\mid Y),$ (19)
@@
A.6 Computational Speed and Memory Analysis
@@
The computational load imposed by any sample that requires additional pass is can be divided into the encoders $f^{(1)},f^{(2)}$ and the fusion network $f_{c}$. The encoders have a computational cost of $cost_{enc}$ per sample, and the fusion network has a cost of $cost_{c}$ per sample. Thus, the overall computational complexity is $\mathcal{O}(M*N*(cost_{enc}+cost_{c}))$, where $M$ is the times we draw noisy samples and $N$ is the batch size. If we now use permutation samples which can be directly drawn from the latent space, the additional computational complexity is reduced to $\mathcal{O}(N*cost_{enc}+M*N*cost_{c}))$. In most state-of-the-art models each modality encoders is significantly larger than the fusion network resulting in $cost_{enc}<<cost_{c}$. In such networks permutations can have almost negligible additional computations. The memory footprint follows a similar pattern.
@@ Appendix B Proof of Supervised Contrastive Loss as Lower Bound
where $\mathcal{P}_{i}=\{p\in\mathbb{I}\mid y_{p}=y_{i}\}$. In supervised contrastive learning, the presence of multiple positive samples turns this into a multi-label problem, unlike traditional NCE methods [39], which typically assume only one positive sample. By taking the version of supervised contrastive learning with the expectation over the positives outside of the log, we can interpret each classification as an average of classifiers, with each classifier focusing on identifying one of the positive samples.
For each positive sample $p\sim\mathcal{P}i$, we aim to derive the optimal probability of correctly identifying that point, denoted $d=p$. This is done by sampling the point from the conditional distribution $p(x{2_{p}}\mid x_{1_{i}},y_{i})$ while sampling the remaining points from the proposal distribution $p(x_{2_{l}})$. This approach mirrors the technique used in InfoNCE [39] and leads to the following derivation:
@@
$L_{\text{Con}}^{Opt}$ $=-\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{i\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[\frac{\frac{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}{p(x_{2_{p}})}}{\frac{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}{p(x_{2_{p}})}+\underset{j\in\mathcal{I},j\neq i}{\sum}\frac{p(x_{2_{j}}\mid x_{1_{i}},y_{i})}{p(x_{2_{j}})}}\right]\right]$ (23)
$=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{i\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[1+\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}\underset{j\in\mathcal{I},j\neq i}{\sum}\frac{p(x_{2_{j}}\mid x_{1_{i}},y_{i})}{p(x_{2_{j}})}\right]\right]$ (24)
$=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{i\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[1+\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}(N-1)\underset{j\in\mathcal{I}}{\mathbb{E}}\frac{p(x_{2_{j}}\mid x_{1_{i}},y_{i})}{p(x_{2_{j}})}\right]\right]$ (25)
$=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{i\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[1+\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}(N-1)\right]\right]$ (26)
$\geq\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{i\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}N\right]\right]$ (27)
@@
Appendix C Additional Ablation Studies
C.1 Comparing Perturbation methods
In this work, we explored three types of perturbation methods to analyze their impact on the performance of MCR: noisy perturbations, zero-masking, and permutations. Each method was applied in different spaces (input space and latent space) or within the batch structure to determine how effectively MCR can leverage these perturbations to enhance multimodal learning.
Table 4 presents the results of an ablation study comparing the performance of MCR under these perturbation techniques across multiple datasets: CREMA-D, AVE, UCF, MOSEI, and MOSI. The methods include:
  1. 1.
    Noise in the Input Space: Adding noise directly to the input features of each modality, simulating realistic data corruption. For its implementation we follow Gat et al. [10].
  2. 2.
    Shapley Input-Space Perturbations: Following the approach of Li et al. [28], Shapley zero-induced values are used to determine the importance of input modalities.
  3. 3.
    Noise in the Latent Space: Applying noise to the latent representations, encouraging robustness at the feature extraction level.
  4. 4.
    Within-Batch Permutations in the Latent Space: Permuting data points within the batch to disrupt alignment.
Table 4: Ablation study comparing different perturbation methods for MCR across multiple datasets. The table shows the performance of MCR when combined with various perturbation techniques, including noise in the input and latent space, Shapley values in the input space, and within-batch permutations.
Datasets
Method CREMA-D AVE UCF MOSEI MOSI
MCR with Noise Input-Space 75.3$\pm$2.9 72.1$\pm$1.1 54.6$\pm$0.8 80.9 76.1
MCR with Shapley Input-Space 73.6$\pm$1.5 72.6$\pm$0.9 55.5$\pm$0.6 81.2 76.2
MCR with Noise Latent-Space 73.6$\pm$1.1 72.6$\pm$0.4 54.5$\pm$0.7 81.3 75.5
MCR with Permutations 76.1$\pm$1.1 73.3$\pm$0.5 55.2$\pm$1.8 81.2 76.2

### v3

@@ A.7 MCR on M modalities
$I(X_{1},..,X_{M};Y)=\hskip-28.45274pt\sum_{\mathcal{S}\subseteq\{X_{1},..,X_{M}\},\mathcal{S}\neq\emptyset}\hskip-28.45274ptI(\mathcal{S};Y\mid\{X_{1},\dots,X_{M}\}\setminus\mathcal{S})+I(X_{1},..,X_{M})-I(X_{1},..,X_{M}\mid Y),$ (21)
@@
A.8 Ablation Study - Game Strategies
We perform an ablation study to test our hypothesis that framing multimodal competition regularization as a game benefits the model by avoiding destructive loss interactions in each backbone encoder. Table 2 compares among the three strategies: Collaborative, Independent, and Greedy. The results show that, across the models allowing backbone encoders to maximize their own $\operatorname{CMI}$ term and concurrently minimizing the others (Greedy strategy) consistently yields the best performance. This result demonstrates that framing multimodal models as competing modalities using game-theoretic principles in the loss terms can be beneficial in balancing these loss terms.
Table 2: Ablation Study: Game Strategies – Comparison of model accuracy on CREMA-D dataset for different game strategies: Collaborative, Independent, and Greedy, using both ResNet and Conformer backbones.
Collaborative Independent Greedy
Dataset/Model Setting $\min\limits_{\theta}\mathcal{L}_{\operatorname{MIPD}}$ $\min\limits_{\theta_{i}}\mathcal{L}_{\operatorname{MIPD}_{X_{i}}}$ $\min\limits_{\theta_{i}}\max\limits_{\theta_{\neg i}}\mathcal{L}_{\operatorname{MIPD}_{X_{i}}}$
[…]
A.9 Ablation Study - Loss Components
Table 3 presents the model’s performance comparison when different loss components are applied. The models utilize pretrained initialization: ResNet with unimodal pretraining and Conformer with SSL. Two key observations can be made:
  1. 1.
    The concurrent exploitation of both $\mathcal{L}_{\operatorname{MIPD}}$ and $\mathcal{L}_{\operatorname{Con}}$ is yielding consistent improvement. Exploiting them separately leads to smaller improvement for $\mathcal{L}_{\operatorname{Con}}$ and even to a decline for $\mathcal{L}_{\operatorname{MIPD}}$. suggests that alignment in the latent space between the modalities is necessary for the permutations to be effective.
  2. 2.
    The $\mathcal{L}_{\operatorname{CEB}}$ term, which penalizes task-irrelevant information, improves the Conformer model’s performance, likely due to its pretraining on large, unlabelled datasets that introduce such irrelevant information. In contrast, for the ResNet model, where pretraining already focuses on task-related information, the $\mathcal{L}_{\operatorname{CEB}}$ term does not provide additional benefits.
Table 3: Ablation Study: Regularization Components – Accuracy (%) of ResNet and Conformer models across datasets with different combinations of $\operatorname{MCR}$ components: $\mathcal{L}_{\operatorname{MIPD}}$, $\mathcal{L}_{\operatorname{Con}}$, and $\mathcal{L}_{\operatorname{CEB}}$. Results indicate that combining $\mathcal{L}_{\operatorname{MIPD}}$ and $\mathcal{L}_{\operatorname{Con}}$ is crucial for improvement, while $\mathcal{L}_{\operatorname{CEB}}$ does not benefit all models.
$\operatorname{MCR}$ Components ResNet Conformer SwinTF
$\mathcal{L}_{\operatorname{MIPD}}$ $\mathcal{L}_{\operatorname{Con}}$ $\mathcal{L}_{\operatorname{CEB}}$ CREMA-D AVE UCF CREMA-D AVE Sth-Sth
[…]
A.10 Ablation Study - Perturbation methods
To estimate the importance of a modality, we perturb one modality while keeping the other fixed and observe the change in the model’s output. Several prior methods have proposed ways to do this, but each comes with trade-offs. Additive noise has been used to maximize output variance as a proxy for functional entropy [18], though this increases sensitivity to noise, conflicting with goals such as smoothness and robustness [66, 67]. Task-specific augmentations [19, 41] rely on handcrafted strategies that may not generalize across domains or modalities. Zero-masking strategies used for approximating Shapley values [21] are theoretically grounded but often unreliable in high-dimensional settings [68] and require multiple forward passes, increasing computational and memory demands.
Based on these previous works we explore three types of perturbation methods to analyze their impact on the performance of $\operatorname{MCR}$: noisy perturbations, zero-masking, and permutations. Each method was applied in different spaces (input space and latent space) or within the batch structure to determine how effectively $\operatorname{MCR}$ can leverage these perturbations to enhance multimodal learning. In Table 4, we summarize the different approaches we examine and in Table 5 we present the results of an ablation study comparing the performance of $\operatorname{MCR}$ under these perturbation techniques across multiple datasets: CREMA-D, AVE, UCF, MOSEI, and MOSI.
Table 4: Overview of the approaches examined for the permutation methods.
Noise in the Input Space Adding noise directly to the input features of each modality, simulating
realistic data corruption. For its implementation we follow [18].
Shapley Input-Space Perturbations Following the approach of [21], Shapley zero-induced values
are used to determine the importance of input modalities.
Noise in the Latent Space Applying noise to the latent representations and encouraging robustness
at the feature extraction level.
Zeros in the Latent Space Zero-masking latent representations to disrupt one modality.
Within-Batch Permutations in the Latent Space Permuting data points within the batch to disrupt alignment.
Table 5: Ablation study comparing different perturbation methods for $\operatorname{MCR}$ across multiple datasets. The table shows the performance of $\operatorname{MCR}$when combined with various perturbation techniques, including input or latent space noise, Shapley values in the input space, and within-batch permutations.
Method CREMA-D AVE UCF MOSEI MOSI
[…]
The semantic meaning of this perturbation depends on whether the permuted sample shares the same label as the original. If the labels match, the perturbation is semantically valid and can be seen as an implicit augmentation. In this case, a large output change indicates that the model may be relying on spurious or unstable features within the modality. If the labels do not match, the resulting input is semantically inconsistent and can be interpreted as out-of-distribution. If the model’s output is insensitive to such a perturbation, this suggests that the modality is being ignored. Conversely, a sensitive reaction may indicate an overreliance on features not robust to semantic shifts. Therefore each category of permuted samples contributes differently to the final output. In practice, we apply both semantically consistent and inconsistent permutations during training. This choice introduces minimal computational overhead and appears to slightly improve convergence stability.
We note that the gradients $\nabla_{\theta_{1}}\mathcal{L}_{\operatorname{MIPD}_{2}}$ and $\nabla_{\theta_{2}}\mathcal{L}_{\operatorname{MIPD}_{1}}$ can negatively impact model robustness depending on the type of perturbation applied. When perturbations yield out-of-distribution unimodal inputs, such as zero-masking or additive noise, the resulting gradients may encourage the model to learn spurious patterns. In contrast, our main experiments use within-batch permutations, which preserve the in-distribution structure of each modality. Under permutation as the perturbation method, the $\mathcal{L}_{\operatorname{MIPD}}$ formulation remains symmetric regardless of which modality is perturbed, and the gradients contribute constructively to learning in all branches.
A.11 Analysis of Multimodal Error
We extend the error analysis from Section 4.3 by comparing unimodal and multimodal predictions in Figure 6. The results echo the pattern seen in CREMA-D (Figure 4): $\operatorname{MCR}$ excels when at least one unimodal model predicts correctly, but still trails MLB and AGM when all unimodal models fail. An exception is the MOSI dataset, where $\operatorname{MCR}$ performs well in synergy, even with three modalities.
Figure 6: Error comparison matrices across pairs of datasets and models, comparing unimodal predictions with multimodal models trained using various methods, including Ensemble, Joint Training, AGM, MLB, and $\operatorname{MCR}$ for the datasets AVE, UCF, MOSI, MOSEI and Something-Something (Sth-Sth). Each column of the confusion matrix represents cases where both unimodal predictions are incorrect, where only one is correct, and where both are correct. The results highlight that $\operatorname{MCR}$ consistently performs well in cases where at least one unimodal prediction is correct. Additionally, MLB and AGM in many instances outperform $\operatorname{MCR}$ in discovering synergetic information, which refers to the "Both/All Wrong" column, highlighting a current limitation of $\operatorname{MCR}$.
A.12 Statistical Importance
We assess the statistical significance of the performance differences between our method and each baseline using the Wilcoxon Signed-Rank Test, applied to per-dataset average results. To control for multiple comparisons, we apply the Holm correction to the resulting $p$-values. We consider results significant if the adjusted $p$-value is below $\alpha=0.05$. Full results are shown in Table 6.
Table 6: Wilcoxon Signed-Rank Test results (two-sided) comparing our method to each baseline across datasets. Statistically significant comparisons after Holm correction ($\alpha=0.05$) are bolded.
Comparison Raw $p$-value Holm-adjusted $p$-value
$\operatorname{MCR}$ vs Ensemble 0.00195 0.02148
$\operatorname{MCR}$ vs Joint Training 0.00195 0.01074
$\operatorname{MCR}$ vs Multi-Loss 0.00195 0.00716
$\operatorname{MCR}$ vs Uni-Pre Frozen 0.00195 0.00537
$\operatorname{MCR}$ vs Uni-Pre Finetuned 0.00195 0.00430
$\operatorname{MCR}$ vs OGM 0.00781 0.00955
$\operatorname{MCR}$ vs AGM 0.00195 0.00358
$\operatorname{MCR}$ vs MLB 0.00195 0.00307
$\operatorname{MCR}$ vs ReconBoost 0.03125 0.03438
$\operatorname{MCR}$ vs MMPareto 0.00195 0.00269
$\operatorname{MCR}$ vs D&R 0.15625 0.15625
Greedy vs Collaborative 0.00977 0.00977
Greedy vs Independent 0.00195 0.00391
$L_{MIPD}$ + $L_{Con}$ + $L_{CEB}$ vs No Reg 0.03125 0.12500
$L_{MIPD}$ + $L_{Con}$ + $L_{CEB}$ vs $L_{MIPD}$ 0.04311 0.08623
$L_{MIPD}$ + $L_{Con}$ + $L_{CEB}$ vs $L_{Con}$ 0.06789 0.09052
$L_{MIPD}$ + $L_{Con}$ + $L_{CEB}$ vs $L_{MIPD}$ + $L_{Con}$ 0.84375 0.84375
Dynamics of MIPD Components
To illustrate the learning dynamics of the $\operatorname{MCR}$ regularizer, we can analyze the evolution of its core loss components during training. Figure 7 plots the two $\operatorname{MIPD}$ terms, corresponding to the video and text modalities, from a training run on the MOSI V-T dataset.
These $\operatorname{MIPD}$ terms, which serve as proxies for the unique contribution of each modality, exhibit a dynamic, alternating behavior. The fluctuations reflect shifts in which modality is more influential on the fused output at different stages of training. This plot visualizes $\operatorname{MCR}$’s mechanism for actively balancing modality importance, preventing one from consistently dominating the other. It is important to note that while this alternating pattern is the desired behavior, its specific form and prominence can vary across different training runs and datasets. Therefore, this figure is presented as a clear, illustrative example of the dynamic interplay $\operatorname{MCR}$ encourages. This prevents static dominance by a single modality, which is key to mitigating modality competition.
Figure 7: The plots show the two $\operatorname{MIPD}$ losses ($\operatorname{MIPD}_{1}$ and $\operatorname{MIPD}_{2}$) for each modality, with (left) and without (right) regularization. With regularization, the losses exhibit an alternating pattern, demonstrating $\operatorname{MCR}$’s ability to dynamically explore multiple modality contribution combinations and prevent collapse into a single modality. Without regularization, the text modality initially contributes more strongly, but the video modality eventually outperforms and dominates.
A.13 Reproduction Challenges for MLA
[…]
A.14 Computational Speed and Memory Analysis of Perturbations
@@
The computational load imposed by any sample that requires an additional pass can be divided into the encoders $f_{1},f_{2}$ and the fusion network $f_{c}$. The encoders have a computational cost of $cost_{enc}$ per sample, and the fusion network has a cost of $cost_{c}$ per sample. Thus, the total computational complexity is $\mathcal{O}((M+1)*N*(cost_{enc}+cost_{c}))$, where $M$ is the times we draw noisy samples to which we add the non-perturbed batch and $N$ is the batch size. If we now use permutation samples that can be directly drawn from the latent space, the additional computational complexity is reduced to $\mathcal{O}(N*cost_{enc}+(M+1)*N*cost_{c})$. Compared to the necessity of a supervised forward pass this translates to an addition of $\mathcal{O}(M*N*cost_{c})$ computational burden. In most state-of-the-art models, each modality encoder is significantly larger than the fusion network, resulting in $cost_{enc}>>cost_{c}$. In such networks, permutations can have almost negligible additional computations. The memory footprint follows a similar pattern.
@@ Appendix B Proof of Supervised Contrastive Loss as Lower Bound
where $\mathcal{P}_{i}=\{p\in\mathbb{I}\mid y_{p}=y_{i}\}$. In supervised contrastive learning, the presence of multiple positive samples turns this into a multi-label problem, unlike traditional noise contrastive estimation (NCE) methods [43], which typically assume only one positive sample. By taking the version of supervised contrastive learning with the expectation over the positives outside of the log, we can interpret each classification as an average of classifiers, with each classifier focusing on identifying one of the positive samples.
For each positive sample $p\sim\mathcal{P}_{i}$, we aim to derive the optimal probability of correctly identifying that point, denoted $d=p$. This is done by sampling the point from the conditional distribution $p(x_{2_{p}}\mid x_{1_{i}},y_{i})$ while sampling the remaining points from the proposal distribution $p(x_{2_{l}})$. This approach mirrors the technique used in InfoNCE [43] and leads to the following derivation:
@@
$L_{\text{Con}}^{Opt}$ $=-\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{p\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[\frac{\frac{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}{p(x_{2_{p}})}}{\frac{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}{p(x_{2_{p}})}+\underset{j\in\mathcal{I},j\neq i}{\sum}\frac{p(x_{2_{j}}\mid x_{1_{i}},y_{i})}{p(x_{2_{j}})}}\right]\right]$ (25)
$=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{p\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[1+\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}\underset{j\in\mathcal{I},j\neq i}{\sum}\frac{p(x_{2_{j}}\mid x_{1_{i}},y_{i})}{p(x_{2_{j}})}\right]\right]$ (26)
$=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{p\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[1+\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}(N-1)\underset{j\in\mathcal{I}}{\mathbb{E}}\frac{p(x_{2_{j}}\mid x_{1_{i}},y_{i})}{p(x_{2_{j}})}\right]\right]$ (27)
$=\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{p\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[1+\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}(N-1)\right]\right]$ (28)
$\geq\underset{i\sim\mathcal{I}}{\mathbb{E}}\left[\underset{p\sim\mathcal{P}_{i}}{\mathbb{E}}\log\left[\frac{p(x_{2_{p}})}{p(x_{2_{p}}\mid x_{1_{i}},y_{i})}N\right]\right]$ (29)